# ForAug: Recombining Foregrounds and Backgrounds to Improve Vision Transformer Training with Bias Mitigation

### v1

@@ Abstract
Transformers, particularly Vision Transformers (ViTs), have achieved state-of-the-art performance in large-scale image classification. However, they often require large amounts of data and can exhibit biases that limit their robustness and generalizability. This paper introduces ForAug, a novel data augmentation scheme that addresses these challenges and explicitly includes inductive biases, which commonly are part of the neural network architecture, into the training data. ForAug is constructed by using pretrained foundation models to separate and recombine foreground objects with different backgrounds, enabling fine-grained control over image composition during training. It thus increases the data diversity and effective number of training samples. We demonstrate that training on ForNet, the application of ForAug to ImageNet, significantly improves the accuracy of ViTs and other architectures by up to 4.5 percentage points (p.p.) on ImageNet and 7.3 p.p. on downstream tasks. Importantly, ForAug enables novel ways of analyzing model behavior and quantifying biases. Namely, we introduce metrics for background robustness, foreground focus, center bias, and size bias and show that training on ForNet substantially reduces these biases compared to training on ImageNet. In summary, ForAug provides a valuable tool for analyzing and mitigating biases, enabling the development of more robust and reliable computer vision models. Our code and dataset are publicly available at https://github.com/tobna/ForAug.
@@ 1 Introduction
Figure 1: Comparison of ForNet and ImageNet. ForNet recombines foreground objects with different backgrounds each epoch, thus creating a more diverse training set. We still apply traditional data augmentation afterwards.
@@
Image classification, a fundamental task in computer vision (CV), involves assigning a label to an image from a predefined set of categories. This seemingly simple task underpins a wide range of applications, including medical diagnosis [39, 50], autonomous driving [52], and object recognition [3, 17, 15]. Furthermore, image classification is used for large-scale pretraining of vision models [10, 31, 47] and to judge the progress of the field of CV [22, 36]. The advent of large-scale datasets, particularly ImageNet [8], containing millions of labeled images across thousands of categories, has been instrumental in driving significant progress in this field. ImageNet served as a catalyst for the rise of large-scale CV models [25, 16] and remains the most important CV benchmark for more than a decade [25, 48, 54, 16].
While traditionally, convolutional neural networks (CNNs) have been the go-to architecture for image classification, Transformers [49], particularly the Vision Transformer (ViT) [10], have emerged as a powerful alternative. These attention-based models have demonstrated superior performance in various vision tasks, including image classification [54, 58, 3, 63, 51].
@@
Data augmentation is a key technique for training image classification models. Traditional data augmentation methods, such as random cropping, flipping, and color jittering, are commonly employed to increase the diversity of the training data and improve the model’s performance [56, 42]. These basic transformations, originally designed for CNNs, change the input images in a way that preserves their semantic meaning [1]. However, the architectural differences of CNNs and Transformers suggest that the latter might benefit from different data augmentation strategies. In particular, the Transformers self-attention mechanism is not translation equivariant [38, 9], meaning that the model does not inherently understand the spatial relationships between pixels.
@@
Inspired by this inductive bias of CNNs, that is not inherent to ViTs, we propose ForAug, a novel data augmentation scheme for image classification which makes the translation equivariance of CNNs explicit in the training data by recombining foreground objects at varying positions with different backgrounds. Applying ForAug to ImageNet gives rise to ForNet, a novel dataset that enables this data augmentation with with fine-grained control over the image composition. Recognizing that Transformers need to learn the spatial relationships from data, since they are not inherently translation invariant, and in general are usually trained on larger datasets [24], we separate the foreground objects in ImageNet from their backgrounds, using an open-world object detector [37], and fill in the background in a plausible way using an object removal model [43, 45]. This allows us to recombine any foreground object with any background on the fly, creating a highly diverse training set. During recombination, we can control important parameters, like the size and position of the foreground object, to help the model learn the spatial invariances necessary for image classification. We show that training on ForNet instead of ImageNet increases the model accuracy of Transformers by up to 4.5 p.p. on ImageNet and an up to $39.3\%$ reduction in error rate on downstream tasks.
@@
Additionally, ForAug is a useful tool for analyzing model behavior and biases, when used during the evaluation phase. We utilize our control over the image distribution to quantify a model’s background robustness (by varying the choice of background), foreground focus (by leveraging our knowledge about the placement of the foreground object), center bias (by controlling the object’s position), and size bias (by controlling object size). These analyses provide insights into model behavior and biases, which is crucial for model deployment and future robustness optimizations. We show that training on ForNet, instead of ImageNet, significantly reduces all of these biases, completely removing the models’ dependence on the background distribution. We make our code for ForAug and the ForNet-dataset publicly available [^1] https://github.com/tobna/ForAug to facilitate further research.
@@
  • We propose ForAug, a novel data augmentation scheme, that recombines objects and backgrounds to train Transformers for image classification.
  • We show that training on ForNet, the ImageNet instantiation of ForAug, leads to 4.5 p.p. improved accuracy on ImageNet and 7.3 p.p. on downstream tasks.
  • We propose novel ForAug-based metrics to analyze and quantify fine-grained biases trained models: Background Robustness, Foreground Focus, Center Bias, and Size Bias. Training on ForNet, instead of ImageNet, significantly reduces these biases.
@@ 2 Related Work
Data augmentation is a crucial technique for improving the performance and generalization of image classification models. Traditional augmentation strategies rely on simple geometric or color-space transformations like cropping, flipping, roatation, blurring, color jittering, or random erasing [62] to increase the diversity of the training data without changing their semantic meaning. With the advent of Transformers, new data augmentation operations like PatchDropout [30] have been proposed. Other transformations like Mixup [61], CutMix [59], or random cropping and patching [46] combine multiple input images. These simple transformations are usually bundled to form more complex augmentation policies like AutoAugment [5] and RandAugment [6], which automatically search for optimal augmentation policies or 3-augment [48] which is optimized to train a ViT. For a general overview of data augmentation techniques for image classification, we refer to [42, 56].
We build upon these general augmentation techniques by introducing a novel approach to explicitly separate and recombine foregrounds and backgrounds for image classification. Our approach is used in tandem with traditional data augmentation techniques to improve model performance and reduce biases.
@@
The copy-paste augmentation [14], which is used for object detection [41, 14] and instance segmentation [53, 28], involves copying segmented objects from one image and pasting them onto another. While typically human-annotated segmentation masks are used to extract the foreground objects, other foregound sources have been explored, like 3D models [19] and pretrained object-detection models for use on objects on white background [11] or synthetic images [12]. DeePaste [53] focuses on using inpainting for a more seamless integration of the pasted object.
Unlike these methods, ForNet focuses on image classification. While for detection and segmentation, objects are pasted onto another image (with a different foreground) or on available or rendered background images of the target scene, we extract foreground objects and fill in the resulting holes in the background in a semantically neutral way. This way, we can recombine any foreground object with a large variety of neutral backgrounds from natural images, enabling a controlled and diverse manipulation of image composition.
@@
Evaluating model robustness to various image variations is critical for understanding and improving model generalization. Datasets like ImageNet-C [18] and ImageNet-P [18] introduce common corruptions and perturbations. ImageNet-E [27] evaluates model robustness against a collection of distribution shifts. Other datasets, such as ImageNet-D [60], focus on varying background, texture, and material, but rely on synthetic data. Stylized ImageNet [13] investigates the impact of texture changes. ImageNet-9 [55] explores background variations using segmented images, but the backgrounds are often artificial.
In contrast to these existing datasets, which are used only for evaluation, ForNet provides fine-grained control over foreground object placement, size, and background selection, enabling a precise and comprehensive analysis of specific model biases within the context of a large-scale, real-world image distribution. As ForNet also provides controllable training set generation, it goes beyond simply measuring robustness to actively improving it through training.
3 RecombiNet (Method)
Figure 2: Overview of ForNet. The data creation consists of two stages: (1, offline) Segmentation, where we segment the foreground objects from the background and fill in the background. (2, online) Recombination, where we combine the foreground objects with different backgrounds to create new samples.
We introduce ForAug, a data augmentation scheme designed to enhance Transformer training by explicitly separating and recombining foreground objects and backgrounds. ForAug involves two stages: Segmentation and Recombination, both visualized in Figure 2.
Segmentation
The segmentation stage isolates the foreground objects and their corresponding backgrounds. We then fill in the background in a visually plausible way [43] using a pretrained object-removal model. This stage is computed once offline and the results are stored for the recombination stage.
@@ 3 RecombiNet (Method)
We iteratively extract up to $n$ foreground masks for each dataset-image, using different more and more general prompts based on the more general synsets of WordNet (e.g. “a sorrel, a type of horse”, “a horse, a type of equine”, …). Masks that are very similar, with a pairwise IoU of at least $0.9$, are merged. The output is a set of masks delineating the foreground objects and the backgrounds. We select the best mask per image (according to Equation 1) in a later filtering step, described below.
An inpainting model that is specifically optimized to remove objects from images, such as LaMa [45] or Attentive Eraser [43], is used to inpaint the foreground regions in the backgrounds. To ensure the quality of the foreground and background images (for each dataset-image), we select a foreground/background pair from the $\leq n$ variants we have extracted and infilled in the previous steps. Using an ensemble of six ViT, ResNet, and Swin Transformer models pretrained on the original dataset, we select the foreground/background pair that maximizes foreground performance while minimizing the performance on the background and size of the foreground according to:
@@
Here, $E$ is the ensemble of models and $m$ is a pretrained model, $c$ is the correct foreground class, $\mathrm{fg}$, and $\mathrm{bg}$ are the foreground and background and $\operatorname{size}(\cdot)$ is the size in number of pixels. We ran a hyperparameter search using a manually annotated subset of foreground/background variants to find the factors in Equation 1: $\lambda=2$ and $\varepsilon=0.1$. The optimal foreground size of $10\%$ of the full image balances the smallest possible foreground size that encompasses all the respective class information in the image with still conveying the foreground information after pasting it onto another background. This filtering step ensures we segment all the relevant foreground objects.
Finally, we filter out backgrounds that are more than $80\%$ infilled, as these tend to be overly synthetic, plain and don’t carry much information (see Appendix C). We ablate this choice in Section 4.1. In summary, we factorize the dataset into a set of foreground objects with a transparent background and a set of diverse backgrounds per class. The next step is to recombine them as data augmentation before applying common data augmentation operations during training.
Recombination
The recombination stage, which is performed online, combines the foreground objects with different backgrounds to create new training samples. For each object, we follow the pipeline of: Pick an appropriate background, resize it to a fitting size, place it in the background image, smooth the transition edge, and apply other data augmentations.
@@
The selected foreground is resized based on its relative size within its original image and the relative size of the original foreground in the selected background image. The final size is randomly selected from a 30% range around upper and lower limits ($s_{u}$ and $s_{l}$), based on the original sizes:
$s\sim\mathcal{U}\left[(1-0.3)s_{l},(1+0.3)s_{u}\right].$ (2)
To balance the size of the foreground and that of the backgrounds original foreground, the upper and lower limit $s_{u}$ and $s_{l}$ are set to the mean or range of both sizes, depending on the foreground size strategy: mean or range.
The resized foreground is then placed at a random position within the background image. This position is sampled from a generalization of the Bates distribution [2] with parameter $\eta\in\mathbb{N}$, visualized in Figure 3. We choose the bates distribution, as it presents an easy way to sample from a bounded domain with just one hyperparameter that controls the concentration of the distribution. $\eta=1$ corresponds to the uniform distribution; $\eta>1$ concentrates the distribution around the center; and for $\eta<-1$, the distribution is concentrated at the borders. To more seamlessly integrate the foreground, we apply a Gaussian blur with ${\sigma\in[\frac{\sigma_{\text{max}}}{10},\sigma_{\text{max}}]}$, inspired by the standard range for the Gaussian blur operation in [48], to the foreground’s alpha-mask.
We can apply standard data augmentation techniques in two modes: Either we apply all augmentations to the recombined image, or we apply the cropping and resizing to the background only and then apply the other augmentations after recombination. The second mode ensures the foreground object remains fully visible, while the first mode mirrors standard data augmentation practices.
We experiment with a constant mixing ratio, or a linear or cosine annealing schedule that increases the amount of images from the original dataset over time. The mixing ratio acts as a probability of selecting an image from the original dataset; otherwise, an image with the same foreground is recombined using ForAug. Thus, we still ensure each foreground is seen once per epoch.
@@ 4 Experiments
We conduct a comprehensive suit of experiments to validate the effectiveness of our approach. We compare training on ForNet, the ImageNet instantiation of ForAug, to training on ImageNet for 7 different models. Furthermore, we assess the impact of using ForNet for pretraining on multiple fine-grained downstream datasets. Additionally, we use ForAug’s control over the image distribution to quantify some model behaviors and biases.
@@ 4.1 Design Choices of ForAug
We start by ablating the design choices of ForAug. For this, we revert to TinyImageNet [26], a subset of ImageNet containing 200 categories with 500 images each, and TinyForNet, a version of ForAug derived from TinyImageNet. Table 1 presents the results of these ablations.
Dataset Detect. Infill FG. Augmentation BG. BG. edge original TinyImageNet Accuracy
prompt Model size Order strategy pruning smoothing image mixing ViT-Ti [%] ViT-S [%]
@@
[…]
Table 1: Ablation of design decisions of TinyForNet on TinyImageNet and ForNet on ImageNet.
@@
Inpainting. Attentive Eraser [43] produces superior results compared to LaMa [45] (see Appendix B for examples).
@@
Background pruning. When it comes to the choice of backgrounds to use, we test two pruning thresholds ($t_{\text{prune}}$) to exclude backgrounds with excessive inpainting. A threshold of $t_{\text{prune}}=1.0$ means that we use all backgrounds that are not fully infilled. Varying $t_{\text{prune}}$ has minimal impact. Therefore, we choose $t_{\text{prune}}=0.8$ to exclude predominantly artificial backgrounds. Similarly, applying edge smoothing to foreground masks with Gaussian blurring actually hurts performance on TinyForNet, but slightly improves it on ForNet.
Mixing ForNet with the original ImageNet data proves crucial. While constant and linear mixing schedules improve performance over no mixing by $2-3$ p.p. compared to only using TinyForNet, the cosine annealing schedule yields the best results, boosting accuracy by another $0.5-1$ p.p.
Background strategy. Another point is the allowed choice of background image for each foreground object. We compare using the original background, a background from the same class, and any background. These strategies go from low diversity and high shared information content between the foreground and background to high diversity and low shared information content. For ViT-Ti, the latter two strategies perform comparably, while ViT-S benefits from the added diversity of using any background. The same is true when training on the full (ImageNet) version of ForNet.
Training Set/ Bates Parameter TIN TinyForNet
$\eta=-3$ $-2$ $/-1$ 1 $2$ $3$
TinyImageNet 68.9 60.5 60.2 60.8 62.6 63.1

### v2

# ForAug: Mitigating Biases and Improving Vision Transformer Training by Recombining Foregrounds and Backgrounds

@@ Abstract
Transformers, particularly Vision Transformers (ViTs), have achieved state-of-the-art performance in large-scale image classification. However, they often require large amounts of data and can exhibit biases, such as center or size bias, that limit their robustness and generalizability. This paper introduces ForAug, a novel data augmentation operation that addresses these challenges by explicitly imposing invariances into the training data, which are otherwise part of the neural network architecture. ForAug is constructed by using pretrained foundation models to separate and recombine foreground objects with different backgrounds. This recombination step enables us to take fine-grained control over object position and size, as well as background selection. We demonstrate that using ForAug significantly improves the accuracy of ViTs and other architectures by up to 4.5 percentage points (p.p.) on ImageNet, which translates to 7.3 p.p. on downstream tasks. Importantly, ForAug not only improves accuracy but also opens new ways to analyze model behavior and quantify biases. Namely, we introduce metrics for background robustness, foreground focus, center bias, and size bias and show that using ForAug during training substantially reduces these biases. In summary, ForAug provides a valuable tool for analyzing and mitigating biases, enabling the development of more robust and reliable computer vision models. Our code and dataset are publicly available at <url>.
@@ 1 Introduction
Figure 1: Comparison of traditional image classification training and training when using ForAug. ForAug recombines foreground objects with different backgrounds each epoch, thus creating a more diverse training set. We still apply strong traditional data augmentation afterwards.
@@
Image classification, a fundamental task in computer vision (CV), involves assigning labels to images from a set of categories. It underpins a wide range of applications, like medical diagnosis [39, 50], autonomous driving [53], and object recognition [2, 16, 14] and facilitates large-scale pretraining [9, 30, 47], and progress evaluation in CV [22, 36]. The advent of large-scale datasets, particularly ImageNet [7], served as a catalyst for the rise of large-scale CV models [24, 15] and remains the most important CV benchmark for more than a decade [24, 48, 56, 15]. While traditionally, convolutional neural networks (CNNs) have been the go-to architecture in CV, Transformers [49], particularly the Vision Transformer (ViT) [9], have emerged as a powerful alternative and go-to architecture, demonstrating superior performance in various vision tasks, including image classification [56, 60, 2, 65, 52].
@@
Data augmentation is a key technique for training image classification models. Traditional augmentation methods, such as cropping, flipping, or color shifts, are commonly employed to increase data diversity [58, 42], but remain bound to existing image compositions. While these preserve the images’ semantic meaning, their ability to teach spatial invariances is limited. While combinations of these data augmentations are still used today, they originally were proposed to benefit CNNs. However, the architectural differences of CNNs and Transformers suggest that the latter might benefit from different data augmentation strategies. In particular, the self-attention mechanism, unlike a CNN, is not translation equivariant [38, 8], meaning that the model is not designed to understand the spatial relationships between pixels.
@@
Recognizing that Transformers need to learn spatial relationships directly from data, we propose ForAug, a data augmentation method that makes these relationships explicit by recombining foreground objects with diverse backgrounds. Thus, ForAug goes beyond existing image compositions and encodes desired invariances directly into the training data (see Figure 1). Applying ForAug to a dataset like ImageNet is a two-step process: (1) We separate the foreground objects in ImageNet from their backgrounds, using an open-world object detector [37] and fill in the background in a neutral way using an object removal model [43, 45]. (2) This allows us to then recombine any foreground object with any background on the fly, creating a highly diverse training set. By exploiting the control over foreground size and position during recombination, ForAug explicitly teaches spatial invariances that image classification models typically must learn implicitly. We show that using ForAug additionally to strong traditional data augmentation increases the model accuracy of Transformers by up to 4.5 p.p. on ImageNet and reduces the error rate by up to $7.3$ p.p. in downstream tasks.
@@
Beyond training, ForAug becomes a diagnostic tool for analyzing model behavior and biases, when used during evaluation. We utilize our control over the image distribution to measure a model’s background robustness (by varying the choice of background), foreground focus (by leveraging our knowledge about the placement of the foreground object), center bias (by controlling position), and size bias (by controlling size). These analyses provide valuable insights into model behavior and biases, which is crucial for model deployment and future robustness optimizations. We show that training using ForAug significantly reduces all of these biases. We make our code for ForAug and the output of ForAug’s segmentation phase on ImageNet publicly available [^1] Link will go here. to facilitate further research.
@@
  • We propose ForAug, a novel data augmentation scheme, that recombines objects and backgrounds. ForAug allows us to move beyond the (possibly biased) image compositions in the dataset while preserving label integrity.
  • We show that training a standard ViT using ForAug leads to up to 4.5 p.p. improved accuracy on ImageNet-1k and 7.3 p.p. on downstream tasks.
  • We propose novel ForAug-based metrics to analyze and quantify fine-grained biases of trained models: Background Robustness, Foreground Focus, Center Bias, and Size Bias. We show that ForAug significantly reduces these biases by encoding invariance that benefits ViT into the training data.
@@ 2 Related Work
Data augmentation is a crucial technique for improving the performance and generalization of image classification models. Traditional augmentation strategies rely on simple geometric or color-space transformations like cropping, flipping, roatation, blurring, color jittering, or random erasing [64] to increase the diversity of the training data without changing their semantic meaning. With the advent of Vision Transformers, new data augmentation operations like PatchDropout [29] have been proposed. Other transformations like Mixup [63], CutMix [61], or random cropping and patching [46] combine multiple input images. These simple transformations are usually bundled to form more complex augmentation policies like AutoAugment [4] and RandAugment [5], or 3-augment [48] which is optimized to train a ViT. For a general overview of data augmentation techniques for image classification, we refer to Shorten and Khoshgoftaar [42], Xu et al. [58].
We build upon these general augmentations by introducing a novel approach to explicitly separate objects and backgrounds for image classification, allowing us to – unlike these basic transformations – move beyond dataset image compositions. Our approach is used additionally to strong traditional techniques to improve performance and reduce biases.
@@
The copy-paste augmentation [13], which is used only for object detection [41, 13] and instance segmentation [54, 27], involves copying segmented objects from one image and pasting them onto another. While typically human annotated segmentation masks are used to extract the foreground objects, other foregound sources have been explored, like 3D models [18] and pretrained object-detection models for use on objects on white background [10] or synthetic images [11]. [20] apply copy-paste as an alternative to CutMix in image classification, but they do not shift the size or position of the foregrounds and use normal dataset images as backgrounds.
Unlike prior copy-paste methods that overlay objects, ForAug extracts foregrounds and replaces their backgrounds with semantically neutral fills, thereby preserving label integrity while enabling controlled and diverse recombination.
Figure 2: Overview of ForAug. The data creation consists of two stages: Segmentation (offline, Section 3.1), where we segment the foreground objects from the background and fill in the background. Recombination (online, Section 3.2), where we combine the foreground objects with different backgrounds to create new samples. After recombination, we apply strong, commonly used augmentation policies.
@@
Evaluating model robustness to various image variations is critical for understanding and improving model generalization. Datasets like ImageNet-C [17] and ImageNet-P [17] introduce common corruptions and perturbations. ImageNet-E [26] evaluates model robustness against a collection of distribution shifts. Other datasets, such as ImageNet-D [62], focus on varying background, texture, and material, but rely on synthetic data. Stylized ImageNet [12] investigates the impact of texture changes. ImageNet-9 [57] explores background variations using segmented images, but backgrounds are often artificial.
In contrast to these existing datasets, which are used only for evaluation, ForAug provides fine-grained control over foreground object placement, size, and background selection, enabling a precise and comprehensive analysis of specific model biases within the context of a large-scale, real-world image distribution. As ForAug also provides controllable training set generation, it goes beyond simply measuring robustness to actively improving it through training.
3 ForAug (Method)
We introduce ForAug, a data augmentation designed to enhance Transformer training by embedding spatial invariances–which Transformers would otherwise need to learn implicitly–directly into the training data. ForAug comprises two distinct stages: Segmentation and Recombination. Both stages are illustrated in Figure 2.
3.1 Segmentation
The segmentation stage isolates the foreground objects and their corresponding backgrounds. We then fill the background using a pretrained object-removal model, producing visually plausible [43], neutral scenes ready for recombination. This stage is computed once offline and the results are stored for the recombination stage.
@@ 3.1 Segmentation
We iteratively extract $n$ foreground masks for each dataset-image, creating prompts by going one hypernym up the WordNet-tree each step (e.g. “a sorrel, a type of horse”, “a horse, a type of equine”, …). Masks that are very similar, with a pairwise IoU of at least $0.9$, are merged. The output is a set of masks delineating the foreground objects and the backgrounds. We select the best mask per image (according to Equation 1) in a later filtering step, described below.
First, an inpainting model that is specifically optimized to remove objects from images, such as LaMa [45] or Attentive Eraser [43], is used to inpaint the foreground regions in the backgrounds. Then, to ensure the quality of the foregrounds and the neutral background images, we select a foreground/background pair (for each dataset-image) from the $\leq n$ variants we have extracted and infilled in the previous steps. Using an ensemble $E$ of six ViT, ResNet, and Swin Transformer models pretrained on the original dataset, we select the foreground/background pair that maximizes foreground performance while minimizing the performance on the background and size of the foreground. For each model $m\in E$, we predict the score of the ground truth class $c$ on the foreground $\mathrm{fg}$ and background $\mathrm{bg}$ and weigh these with the size $\operatorname{size}(\cdot)$ in number of pixels according to:
@@
We run a hyperparameter search using a manually annotated subset of foreground/background variants to find the factors in Equation 1: $\lambda=2$ and $\varepsilon=0.1$.
Finally, we filter out backgrounds that are largely infilled, as these tend to be overly synthetic and do not carry much information (see the supplementary material). Although the segmentation stage is computational overhead, it is a one-time cost with results that can be reused across experiments (see the supplementary material for details). In summary, we factorize the dataset into a set of foreground objects with a transparent background and a set of diverse backgrounds per class. The next step is to recombine these, before applying other common data augmentation operations during training.
3.2 Recombination
The recombination stage, performed online during training, combines the foreground objects with different backgrounds to create new training samples. For each object, we follow the pipeline of: Pick an appropriate background, resize it to a fitting size, and place it in the background image. Through this step, we expose the model to variations beyond the image compositions of the dataset.
@@ 3.2 Recombination
The selected foreground is resized based on its relative size within its original image and the relative size of the original foreground in the selected background image. The final size is randomly selected from a 30% range around upper and lower limits ($s_{u}$ and $s_{l}$), based on the original sizes. To balance the size of the foreground and that of the backgrounds original foreground, the upper and lower limit $s_{u}$ and $s_{l}$ are set to the mean or range of both sizes, depending on the foreground size strategy: mean or range.
The resized foreground is then placed at a random position within the background image. To more seamlessly integrate the foreground, we apply a Gaussian blur with $\sigma\in[\frac{\sigma_{\text{max}}}{10},\sigma_{\text{max}}]$}, inspired by the standard range for the Gaussian blur operation in [48], to the foreground’s alpha-mask.
We can apply standard data augmentation techniques in two modes: Either we apply all augmentations to the recombined image, or we apply the cropping and resizing to the background only and then apply the other augmentations after recombination. The first mode mirrors standard augmentation practice, whereas the second one ensures the foreground object remains fully visible.
We experiment with a constant mixing ratio, or a linear or cosine annealing schedule that increases the amount of images from the original dataset over time. The mixing ratio acts as a probability of selecting an image from the original dataset; otherwise, an image with the same foreground is recombined using ForAug, ensuring each object is seen once per epoch. The recombination stage is designed to be parallelized on the CPU during training and thus does not impact training time (see supplementary material for details).
@@ 4 Experiments
We conduct a comprehensive suit of experiments to validate the effectiveness of our approach, comparing ImageNet-training with and without ForAug for 10 different models. Furthermore, we assess the impact of using ForAug for pretraining on multiple fine-grained downstream datasets. Finally, we exploit ForAug’s control over the image distribution to quantify model behaviors and biases. We always report the mean and standard deviation of three independent training runs.
@@ 4.1 Design Choices of ForAug
We start by ablating the design choices of ForAug on TinyImageNet [25], a subset of ImageNet containing 200 categories with 500 images each. Table 1 presents ablations for segmentation and Table 2 for recombination.
Table 1: Ablation of the design decisions in the segmentation phase of ForAug on TinyImageNet. The first line is our baseline, while the other lines are using ForAug. We use basic settings with the same background strategy during recombination for this experiment.
Detect. Prompt Infill Model TinyImageNet Accuracy [%]
ViT-Ti ViT-S
TinyImageNet $66.1\pm 0.5$ $68.3\pm 0.7$
specific LaMa [45] $65.5\pm 0.4$ $71.2\pm 0.5$
general LaMa [45] $66.4\pm 0.6$ $72.9\pm 0.6$
general Att. Eraser [43] $67.5\pm 1.2$ $72.4\pm 0.5$
Table 2: Ablation of the recombination phase of ForAug on TinyImageNet (top) and ImageNet (bottom). The first experiments use the initial segmentation settings with LaMa [45].
FG.
size
Augment.
Order
BG
Strat.
BG.
Prune
Original
Mixing
Edge
Smooth.
Accuracy [%]
ViT-Ti ViT-S
@@
Inpainting. Among inpainting models, Attentive Eraser [43] produces slightly better results compared to LaMa [45] (+0.5+0.5 p.p. on average). For inpainting examples, see the supplementary material.
@@
Background pruning. When it comes to the backgrounds to use, we test different pruning thresholds ($t_{\text{prune}}$) to exclude backgrounds with large inpainting. A threshold of $t_{\text{prune}}=1.0$ means that we use all backgrounds that are not fully infilled. Varying $t_{\text{prune}}$ has minimal impact. We choose $t_{\text{prune}}=0.8$ to exclude predominantly artificial backgrounds.
Mixing ForAug-augmented samples with the original ImageNet data proves crucial. While constant and linear mixing schedules improve performance over no mixing by 232-3 p.p. compared to only augmented samples, the cosine annealing schedule proves optimal, boosting accuracy by 343-4 $p.$
Edge smoothing. We evaluate the impact of using Gaussian blurring to smooth the edges of the foreground masks. For larger models, this gives us a slight performance boost on the full ImageNet (second to last line in Table 2).
Background strategy. Another point is the allowed choice of background image for each foreground object. We compare using the original background, a background from the same class, and any background. These strategies go from low diversity and high shared information content between the foreground and background to high diversity and low shared information content. For ViT-Ti, the latter two strategies perform comparably, while ViT-S benefits from the added diversity of using any background. The same is true when training on the full ImageNet.
Table 3: Accuracy of ViT-S on TinyImageNet (TIN) in percent using ForAug with different foreground position distributions by varying the Bates parameter $\eta$. The best performance is achieved when using the uniform distribution ($\eta=1$) for training.
Bates Parameter during training TIN w/o ForAug TIN w/ ForAug
$\eta=-3$ 2-2 1/11/-1 22 33
Baseline 68.9 60.5 60.2 60.8 62.6 63.1